# RobGC: Towards Robust Graph Condensation

### v1

@@ IV-F Ablation Study (Q6)
Figure 4: The hyper-parameter sensitivity analysis for the number of selected nearest neighbors and hops. The number of selected nearest neighbors controls the graph sparsity and the number of hops decides the candidate nodes.
@@ IV-G Analysis on Optimized Graph (Q7)
To assess the impact of the structure optimization, we compare several properties between optimized graphs and original noisy graphs in Table VIII. Here, we employ the widely used edge homophily measurement [34], which is defined as the fraction of edges in the graph which connects nodes that have the same class label. The comparison reveals an overall improvement in homophily rates on optimized graphs, with a notable observation that the enhancement in test graphs is not as pronounced as in training graphs. The potential rationale is that the thresholds are optimized according to the training graph and directly applied to test data for efficiency. Regarding graph sparsity, we observe an increase in the number of edges within the Cora, Citeseer, and Pubmed datasets. This finding aligns with insights from our ablation study in Table VII, indicating that edge addition plays a more crucial role than deletion in these specific graphs. To obtain the optimal results across different datasets, it’s crucial to implement both edge deletion and addition strategies in the structure optimization.
@@ IV-H Hyper-parameter Sensitivity Analysis (Q8)
In this section, we evaluate the number of selected nearest neighbors and hops for neighbor candidates in the locality-based edge addition. In Fig. 4, we present the performances w.r.t. different values of the number of selected nearest neighbors and hops for neighbor candidates on Cora and Citeseer, respectively. Similar trends were noted across other datasets. For the number of selected nearest neighbors, we observe that choosing 2 to 3 nearest neighbors yields the best performance. Higher values may include an excessive number of neighbors, resulting in a denser adjacency matrix that could potentially introduce more noisy edges. On the other hand, the performance shows relative insensitivity to the number of hops for neighbor candidates. While a higher number of hops can encompass more potentially effective neighbors, it also significantly increases the number of candidates and computational complexity. Therefore, we select the number of hop as 3 in the experimental results to maintain a balance between comprehensive neighbor selection and manageable computation demands.
@@ V-A Graph Condensation
Graph condensation [14, 42, 43] is designed to reduce GNN training costs through a data-centric perspective and has been wildly used in various applications due to its excellent graph compression performance, including inference acceleration [44], continual learning [45], hyper-parameter/neural architecture search [46] and federated learning [12].
Existing research efforts can be broadly categorized into four classes based on the primary criteria of GC: effective GC, generalized GC, fair GC, and efficient GC. Each kind of method try to optimize and enhance the condensation procedure from different aspects. Firstly, effective GC prioritize improving the accuracy of GNNs trained on condensed graphs. These approaches often employ advanced optimization strategies to enhance the optimization results, such as trajectory matching[33], kernel ridge regression [47], and graph neural tangent kernels[33]. These strategies are designed to ensure that the condensed graph retains essential structural and feature information critical for high GNN performance. Generalized GC aims to improve the overall performance of different GNN architectures trained on the condensed graph. The primary challenge of generalization is to minimize the loss of information in the structural during condensation.For example, SGDD [48] employs the Laplacian energy distribution to analyze the structural properties of the graph, quantifying the spectral shift between condensed and training graphs. GCEM [32] circumvents the conventional relay GNN and directly generates the eigenbasis for the condensed graph, eliminating the spectrum bias inherent in relay GNNs. Fair GC focuses on the disparity in fairness between models trained on condensed graphs and those trained on training graphs [49], aiming to prevent bias amplification in the condensed graph and ensure equitable model performance. Finally, efficient GC [31] concentrates on the speed of condensed graph generation and the method’s feasibility for time-sensitive scenarios. These methods are developed to accelerate the condensation process across all components within GC, including: graph encoding, optimization, and graph generation. By enhancing the efficiency of each stage, efficient GC methods can significantly reduce the time required for the entire GC process.
@@ V-B Graph Structure Learning
The performance of GNNs is significantly influenced by the quality of the graph structures, as these structures play a crucial role in learning informative node embeddings. To enhance the robustness of GNN models, graph structure learning has been introduced to jointly optimize the graph structure while learning corresponding node representations. The essence of graph structure learning lies in modeling the optimal graph structure, which can be categorized into three classes: metric learning, probabilistic modeling, and direct optimization [28].
Specifically, metric learning approaches model edge weights as the distance between each pair of node representations and introduces numerous metric functions like Gaussian kernel [50], inner product [51], and cosine similarity [52]. Each metric offers a distinct approach to defining the similarity between nodes, thereby guiding the formation of edges and influencing the graph structure optimization process. Probabilistic modeling approaches [53, 54] assume that the graph is generated by a specific distribution. Consequently, these approaches use a trainable model to define the edge probability distribution and apply re-parameterization techniques to learn the discrete graph structure. Direct optimization approaches [55, 56] consider the optimized graph adjacency matrix as learnable parameters and optimize the adjacency matrix toward achieving a feasible graph structure. Therefore, they employ regularization techniques informed by prior knowledge of graphs, emphasizing aspects like feature smoothness [57], sparsity [58], and the low-rank nature [27] of the graph structure. Nevertheless, these methods demand intensive computations for every node pair in the graph, leading to a computational complexity of $\mathcal{O}(N^{2})$ for a graph with $N$ nodes. To address this challenge, recent attempts like SUBLIME [35] and NodeFormer [59] have adopted approximation techniques and transformer architectures to reduce the complexity. However, the optimization of the graph structure within these methods is tightly entangled with the specific model. As a result, any change to the GNN model requires retraining the methods from the beginning, introducing a lack of flexibility and additional training time.
@@ VI Conclusion
In this paper, we present robust graph condensation (RobGC), a general approach for GC to extend the robustness and applicability of condensed graphs in noisy scenarios. RobGC incorporates a denoising process to tackle noise present in the training graph and adopts an alternating optimization strategy for condensation and denoising processes, contributing to the mutual promotion of the condensed graph and the training graph. Additionally, RobGC facilitates the inductive test-time graph denoising by leveraging the noise-free condensed graph, preventing the inductive nodes from noise during inference. As a preliminary exploration, RobGC investigates the structural noise in training and test graph. The future work can further explore the robustness of GC with node feature noise and label noise to extend the applicability of GC in a broader array of scenarios.

### v2

@@ IV-F Ablation Study (Q6)
Besides, to verify the convergence of the alternating optimization process, we present the performance of the GCN model trained on both the original and condensed graphs throughout the optimization procedure. The results for all datasets are shown in Fig. 4. We observe that the accuracy of the model trained on the original graph gradually improves and converges as the condensation procedure progresses. This demonstrates that the quality of the condensed graph significantly influences the denoising performance of the original graph. Furthermore, as the quality of the original graph improves, the performance of the condensed graph also exhibits gradual enhancement, highlighting the effectiveness of our proposed GC method.
Figure 4: The accuracy of the model trained on the original and condensed graphs throughout our optimization procedure. The GCond and random noise are applied and the noise level is 100%. The condensation ratios are 5.80%, 3.30%, 0.17%, and 0.50% for datasets, respectively.
Figure 5: The hyper-parameter sensitivity analysis for the number of selected nearest neighbors, hops, and searching candidate on Citeseer dataset.
@@ IV-G Analysis on Optimized Graph (Q7)
To assess the impact of the structure optimization, we compare several properties between optimized graphs and original noisy graphs in Table VIII. Here, we employ the widely used edge homophily measurement [44], which is defined as the fraction of edges in the graph which connects nodes that have the same class label.
The comparison reveals an overall improvement in homophily rates on optimized graphs, with a notable observation that the enhancement in test graphs is not as pronounced as in training graphs. The potential rationale is that the thresholds are optimized according to the training graph and directly applied to test data for efficiency. Specifically, compared with the noisy test graph, the number of edges decreases after edge deletion and then increases by edge addition. Moreover, homophily consistently increases throughout two phases, demonstrating the effectiveness of both unreliable edge deletion and addition phases. Regarding graph sparsity, we observe an increase in the number of edges within the Cora, Citeseer, and Pubmed datasets. This finding aligns with insights from our ablation study in Table VII, indicating that edge addition plays a more crucial role than deletion in these specific graphs. To obtain the optimal results across different datasets, it’s crucial to implement both edge deletion and addition strategies in the structure optimization.
In addition, Fig. 6 presents the t-SNE visualization of node embeddings in the original graph and condensed graph generated by our proposed RobGC. The condensed graph simplifies intra-class distributions while preserving the overall positioning and relationships between classes, ensuring that key inter-class structures are retained. This underscores the representativeness of the condensed node distributions.
Figure 6: The visualization of the original (left) and condensed (right) graphs, where different colors represent distinct classes.
TABLE IX: The performance of our method combined with robust GNNs. The random noise is applied and the noise level is 100%. The condensation ratios are 5.80%, 3.30%, 0.17%, and 0.50% for datasets, respectively.
Cora Citeseer Pubmed Ogbn-arxiv
GCond+GCN 64.7±0.8 66.2±0.1 69.8±0.3 46.4±0.5
GCond+${\rm{STABLE_{M}}}$ 65.5±0.7 67.6±0.2 70.2±0.5 47.5±0.6
Ours+${\rm{STABLE_{M}}}$ 70.9±0.3 73.2±0.1 77.0±0.3 50.1±0.3
GCond+Soft Median 65.8±0.5 67.9±0.4 70.8±0.3 47.9±0.5
Ours+Soft Median 71.2±0.4 73.9±0.3 77.3±0.2 50.6±0.3
@@ IV-H Hyper-parameter Sensitivity Analysis (Q8)
In this section, we evaluate the number of selected nearest neighbors and hops for neighbor candidates in the locality-based edge addition, as well as the number of searching candidates $h$ for optimization. In Fig. 5, we present the performances on Citeseer and similar trends were noted across other datasets. For the number of selected nearest neighbors, we observe that choosing 2 to 3 nearest neighbors yields the best performance. Higher values may include an excessive number of neighbors, resulting in a denser adjacency matrix that could potentially introduce more noisy edges. On the other hand, the performance shows relative insensitivity to the number of hops for neighbor candidates. While a higher number of hops can encompass more potentially effective neighbors, it also significantly increases the number of candidates and computational complexity. For the number of searching candidates, search time increase with larger values of $h$, and accuracy levels off beyond a certain point. Therefore, we select the number of hop as 3 and $h$ as 20 in the experiments to maintain a balance between performance and computation demands.
IV-I Performance for Robust GNNs (Q9)
To further evaluate the generalizability of our method, we evaluate two robust GNNs: ${\rm{STABLE_{M}}}$ [25] and Soft Median [22], on the denoised condensed graphs. Table IX compares data-centric methods for generating condensed graphs (GCond and Ours) and various GNN models (GCN, ${\rm{STABLE_{M}}}$, and Soft Median). The results show that robust GNNs outperform the vanilla GCN model across different datasets. However, their performance is still impacted by the noisy condensed graph. When combined with our RobGC, their performance improves significantly, underscoring the effectiveness of our denoising strategy.
@@ V-A Graph Condensation
Existing graph condensation [12, 45] research can be broadly categorized into 4 classes based on the primary criteria: effective GC, generalized GC, fair GC, and efficient GC. (1) Effective GC prioritizes improving the accuracy of GNNs trained on condensed graphs. These approaches often employ advanced optimization strategies to enhance the optimization results, such as trajectory matching [32], kernel ridge regression [46], and graph neural tangent kernels [32]. Moreover, recent studies explore the advanced techniques to explore the inherent information in the original graph, including self-supervised learning [47], difficult sample mining [33], and self-expressive ability [48]. (2) Generalized GC [49] aims to improve the overall performance of different GNN architectures trained on the condensed graph. The primary challenge of generalization is to minimize the loss of information in the structure during condensation. (3) Fair GC [50] focuses on the disparity in fairness between models trained on condensed graphs and those trained on training graphs, aiming to prevent bias amplification in the condensed graph and ensure equitable model performance. (4) Efficient GC [27] concentrates on the speed of graph generation and the method’s feasibility for time-sensitive scenarios. These methods are developed to accelerate the condensation process across all components in GC, including: graph encoding [31], optimization [51], and graph generation [52].
Besides these methods, another related GC study is MCond [53], which addresses the high latency issue in the same inductive inference setting as ours. It introduces a mapping matrix that connects the inductive graph to the condensed graph for efficient message-passing. Building on this insight, RobGC can be further refined to enable inductive graphs to perform efficient message-passing on noise-free condensed graphs, with significant potential to enhance GC performance in both resource-constrained, and noisy scenarios.
@@ V-B Graph Structure Learning
To enhance the robustness of GNN models, graph structure learning has been introduced to jointly optimize the graph structure while learning corresponding node representations, which can be categorized into three classes: metric learning, probabilistic modeling, and direct optimization.
Specifically, metric learning approaches model edge weights as the distance between each pair of node representations and introduces numerous metric functions like Gaussian kernel [54], inner product [55], and cosine similarity [56]. Each metric offers a distinct approach to defining the similarity between nodes, thereby guiding the formation of edges and influencing the graph structure optimization process. Probabilistic modeling approaches [57, 58] assume that the graph is generated by a specific distribution. Consequently, these approaches use a trainable model to define the edge probability distribution and apply re-parameterization techniques to learn the discrete graph structure. Direct optimization approaches [59, 60] consider the optimized graph adjacency matrix as learnable parameters and optimize the adjacency matrix toward achieving a feasible graph structure. Therefore, they employ regularization techniques informed by prior knowledge of graphs, emphasizing aspects like feature smoothness [61], sparsity [62], and the low-rank nature [40] of the graph structure.
@@ VI Conclusion
In this paper, we present robust graph condensation (RobGC), a general approach for GC to extend the robustness and applicability of condensed graphs in noisy scenarios. RobGC incorporates a denoising process to tackle noise present in the training graph and adopts an alternating optimization strategy for condensation and denoising processes, contributing to the mutual promotion of the condensed graph and the training graph. Additionally, RobGC facilitates the inductive test-time graph denoising by leveraging the noise-free condensed graph, preventing the inductive nodes from noise during inference. Our empirical analysis demonstrates that adversarial attacks pose a substantial threat to the quality of graph structures. Future research could further investigate the advanced relay model in GC, incorporating robust aggregation functions [24] to filter out noise and counteract the effects of adversarial attacks on both features and labels.